OpenStudy (anonymous):

In a circle of radius 10 cm, a sector has an area of 40 sq. cm. What is the degree measure of the arc of the sector?

OpenStudy (saifoo.khan):

\[Area = \frac{\theta}{360} \times \pi r^2\]